# Hypergraph Transformer for Semi-Supervised Classification

### v1

@@ Abstract
Hypergraphs play a pivotal role in the modelling of data featuring higher-order relations involving more than two entities. Hypergraph neural networks emerge as a powerful tool for processing hypergraph-structured data, delivering remarkable performance across various tasks, e.g., hypergraph node classification. However, these models struggle to capture global structural information due to their reliance on local message passing. To address this challenge, we propose a novel hypergraph learning framework, HyperGraph Transformer (HyperGT). HyperGT uses a Transformer-based neural network architecture to effectively consider global correlations among all nodes and hyperedges. To incorporate local structural information, HyperGT has two distinct designs: i) a positional encoding based on the hypergraph incidence matrix, offering valuable insights into node-node and hyperedge-hyperedge interactions; and ii) a hypergraph structure regularization in the loss function, capturing connectivities between nodes and hyperedges. Through these designs, HyperGT achieves comprehensive hypergraph representation learning by effectively incorporating global interactions while preserving local connectivity patterns. Extensive experiments conducted on real-world hypergraph node classification tasks showcase that HyperGT consistently outperforms existing methods, establishing new state-of-the-art benchmarks. Ablation studies affirm the effectiveness of the individual designs of our model.
@@ 4.1 Hypergraph incidence matrix based positional encoding
Theorem 1. Let $\mathbf{p}_{u}$, $\mathbf{p}_{v}$ be the positional encoding of nodes $u$ and $v$, respectively. Then, $\lVert\mathbf{p}_{u}-\mathbf{p}_{v}\rVert_{2}\leq C\sqrt{N_{e}}$,
where $C$ is a constant and $N_{e}$ is the number of hyperedges only with either node $u$ or $v$.
@@
where $\operatorname{\textbf{h}}_{u},\operatorname{\textbf{h}}_{v}$ are the $u$-th and $v$-th row vectors of $\mathbf{H}$, $\lVert\cdot\rVert_{2}$ denotes the 2-norm of a vector or a matrix, and $\sigma_{max}(\cdot)$ denotes the largest singular value of a matrix. From Eq. (2), we know that
the distance between the positional encodings of the two nodes is bounded by the number of hyperedges that only contain either node $u$ or $v$, while the latter represents the similarity between the structural information of node $u$ and $v$. We have similar results for hyperedges as well.
@@ 4.4 Training strategy
Given training labels $\mathbf{Y}_{lab}=\{\mathbf{y}_{u}\}_{u\in\mathcal{V}_{lab}}$, where $\mathcal{V}_{lab}$ denotes the set of labeled nodes, the common practice is to maximize the observed data log-likelihood which yields a supervised loss with all classes:
@@
5 Experiments
5.1 Experimental settings
Implementation details. Our model is built upon the implementation of Nodeformer [25], which leverages a kernerlized gumbel-softmax operator to reduce the attention computation comlexity to linearity. Regarding the hyperparameters of the $\mathcal{O}(N)$ Transformer, we set them to be consistent with NodeFormer. We randomly split the data (node indices) into training/validation/test samples using 50%/25%/25% splitting percentages as in Chien et al. [13]. We report the average prediction accuracy over ten random splits as the evaluation metric. All experiments are conducted on an RTX 3090 utilizing the PyTorch framework. Our code will be available at: https://github.com/zeroxleo/HyperGT.
5.2 Results and analysis
Comparison with SOTA hypergraph neural networks. The results of our experiments on real-world hypergraph node classification benchmarks are presented in Table 2. In comparison to previous hypergraph neural networks, our proposed HyperGT demonstrates superior performance across all datasets. A key distinguishing feature of HyperGT is its incorporation of both global and local hypergraph structural information, setting it apart from prior methods that solely focus on local structural information. These results show the importance of leveraging global information for designing more effective models in processing hypergraph-structured data.
@@ 5.2 Results and analysis
HyperGT(Ours) 95.09 ± 1.16 65.49 ± 5.11 69.83 ± 0.39 74.55 ± 1.99
Ablation studies. We perform a series of ablation studies on the importance of designs in our proposed HyperGT on the walmart dataset. The ablation results are presented in Table 3.
We see that the use of positional encodings for nodes leads to a significant improvement, which shows that the model effectively captures the hypergraph structural information of nodes. Positional encodings of hyperedges also contribute to positive effects. With the addition of regularization, the performance is further improved, which indicates that this regularizer on hypergraph structure helps the Transformer in learning the specific connections between nodes and hyperedges. All of the components designed are helpful to our Transformer architecture for modelling hypergraph data and help to leverage useful information from input hypergraphs.

### v2

@@ Abstract
Hypergraphs play a pivotal role in the modelling of data featuring higher-order relations involving more than two entities. Hypergraph neural networks emerge as a powerful tool for processing hypergraph-structured data, delivering remarkable performance across various tasks, e.g., hyrgraph node classification. However, these models struggle to capture global structural information due to their reliance on local message passing. To address this challenge, we propose a novel hypergraph learning framework, HyperGraph Transformer (HyperGT). HyperGT uses a Transformer-based neural network architecture to effectively consider global correlations among all nodes and hyperedges. To incorporate local structural information, HyperGT has two distinct designs: i) a positional encoding based on the hypergraph incidence matrix, offering valuable insights into node-node and hyperedge-hyperedge interactions; and ii) a hypergraph structure regularization in the loss function, capturing connectivities between nodes and hyperedges. Through these designs, HyperGT achieves comprehensive hypergraph representation learning by effectively incorporating global interactions while preserving local connectivity patterns. Extensive experiments conducted on real-world hypergraph node classification tasks showcase that HyperGT consistently outperforms existing methods, establishing new state-of-the-art benchmarks. Ablation studies affirm the effectiveness of the individual designs of our model.
@@ 4.1 Hypergraph incidence matrix based positional encoding
Theorem 1. Let $\mathbf{p}_{u}$, $\mathbf{p}_{v}$ be the positional encoding of nodes $u$ and $v$, respectively. Then, $\lVert\mathbf{p}_{u}-\mathbf{p}_{v}\rVert_{2}\leq C\sqrt{N_{e}}$, where $C$ is a constant and $N_{e}$ is the number of hyperedges only with either node $u$ or $v$.
@@
where $\operatorname{\textbf{h}}_{u},\operatorname{\textbf{h}}_{v}$ are the $u$-th and $v$-th row vectors of $\mathbf{H}$, $\lVert\cdot\rVert_{2}$ denotes the 2-norm of a vector or a matrix, and $\sigma_{max}(\cdot)$ denotes the largest singular value of a matrix. From Eq. (2), we know that the distance between the positional encodings of the two nodes is bounded by the number of hyperedges that only contain either node $u$ or $v$, while the latter represents the similarity between the structural information of node $u$ and $v$. We have similar results for hyperedges as well.
@@ 4.4 Training strategy
For node classification, given training labels $\mathbf{Y}_{lab}=\{\mathbf{y}_{u}\}_{u\in\mathcal{V}_{lab}}$, where $\mathcal{V}_{lab}$ denotes the set of labeled nodes, we train our model with the cross-entropy loss function:
@@
HyperGT(Ours) 95.23 ± 0.73 65.49 ± 5.11 69.83 ± 0.39 74.55 ± 1.99
5 Experiments
5.1 Experimental settings
Implementation details. Our model is built upon the implementation of Nodeformer [25], which leverages a kernerlized gumbel-softmax operator to reduce the attention computation comlexity to linearity. We identified $\lambda$’s optimal value via grid-search. Regarding the hyperparameters of the $\mathcal{O}(N)$ Transformer, we set them to be consistent with NodeFormer. We randomly split the data (node indices) into training/validation/test samples using 50%/25%/25% splitting percentages as in Chien et al. [13]. We report the average prediction accuracy over ten random splits as the evaluation metric. All experiments are conducted on an RTX 3090 utilizing the PyTorch framework. Our code is available at: https://github.com/zeroxleo/HyperGT.
5.2 Results and analysis
Comparison with SOTA hypergraph neural networks. The results of our experiments on real-world hypergraph node classification benchmarks are presented in Table 2. In comparison to previous hypergraph neural networks, our proposed HyperGT demonstrates superior classification accuracy across all datasets. A key distinguishing feature of HyperGT is its incorporation of both global and local hypergraph structural information, setting it apart from prior methods that solely focus on local structural information. These results show the importance of leveraging global information for designing more effective models in processing hypergraph-structured data. We also conducted experiments on the Walmart dataset to obtain the inference speed of popular HGNNs and HyperGT. Results are presented in Table 4, indicating that HyperGT maintains competitive inference speeds, aligning with its theoretically low linear complexity, while also preserving global attention interactions.
Ablation studies. We perform a series of ablation studies on the importance of designs in our proposed HyperGT on the walmart dataset. The ablation results are presented in Table 3. We see that the use of positional encodings for nodes leads to a significant improvement, which shows that the model effectively captures the hypergraph structural information of nodes. Positional encodings of hyperedges also contribute to positive effects. With the addition of regularization, the performance is further improved, which indicates that this regularizer on hypergraph structure helps the Transformer in learning the specific connections between nodes and hyperedges. All of the components designed are helpful to our Transformer architecture for modelling hypergraph data and help to leverage useful information from input hypergraphs.
@@ 5.2 Results and analysis
Table 4: Inference speeds (ms/run, mean of 1000 runs) on the Walmart dataset. Results show the competitive inference speed of HyperGT.
HGNN HCHA UniGCNII AllDeepset AllSetTransformer HyperGT
20.301 20.942 25.015 32.603 62.788 24.882